Circular No. 6503

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bt IN ANONYMOUS GALAXY P. Garnavich, Harvard-Smithsonian Center for Astrophysics, reports

that P. Berlind, Fred L. Whipple Observatory (FLWO), has discovered a supernova in an elliptical galaxy located at α = 6^{h}51^{m}58^{s}.0, δ = +16^{o}17′43′′ (equinox 2000.0). The supernova was recognized by its spectrum, obtained on Nov. 10.49 UT with the Tillinghast 1.5-m telescope in the course of the ‘Winter Plane’ redshift survey (principal investigator C. Pantoja). The spectrum shows SN 1996bt to be a type-Ia supernova ∼ 10 days after maximum. CCD images obtained by E. Barton with the FLWO 1.2-m telescope on Nov. 10.50 show the supernova to be 1′′.5 west and 1′′.2 north of the galaxy’s center. Accurate photometry is difficult, due to the blending of the supernova and galaxy images, but point-spread-function fitting gives a magnitude of B = 16.5 ± 0.1.

SGR 1806–20 C. Kouveliotou, Universities Space Research Association; J. van Para-

dijs, University of Amsterdam, and University of Alabama in Huntsville (UAH); G. J. Fishman and C. A. Meegan, Marshall Space Flight Center, NASA; S. Dieters and M. S. Briggs, UAH; K. Hurley, University of California, Berkeley; T. Murakami, Institute of Space and Aeronautical Sciences, I. Smith, Rice University; and D. Frail, National Radio Astronomical Observatory, report: “We observed SGR 1806–20 on Nov. 5 and 6 with the Rossi X-Ray Timing Explorer during an active episode detected by BATSE (IAUC 6501). In addition to the large outbursts, also seen with BATSE, we detected numerous small outbursts coming in groups from the source, with peak fluxes ∼ 2 orders of magnitude fainter than those of bursts seen earlier with BATSE. Burst rates as high as one per minute were observed.”

COMET 29P/SCHWASSMANN-WACHMANN 1 I. Ferrin, J. Mateu, and J. Canelon, University of the Andes, Merida,

report that CCD observations of comet 29P made with the 1-m Schmidt telescope of the National Observatory of Venezuela on Nov. 7.37 UT show a sharp nucleus embedded in a coma of diameter ∼ 33′′. The total magnitude was R = 14.5 ± 0.1. Material is being ejected in a position angle of 115^{o} (solar direction 127^{o}). The comet shows a ‘C’-shaped coma, implying that the nucleus is rotating counterclockwise.

1996 November 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green